OpenStudy (anonymous):

Infinite number of solutions! Find a and b. (2a-1)x-3y=5;3x+(b-2)y=3; As the equations have infinite number of solutions, a1/a2=b1/b2=c1/c2. Here, (2a-1)/3=-3/(b-2)=5/3. How to proceed. I am stuck at the cross multiplying and finding out both a and b.

OpenStudy (anonymous):

Wait, are you given the information that the two equations have an infinite number of solutions?

OpenStudy (anonymous):

Yes! @genius12

OpenStudy (anonymous):

What is the a1/a2 = b1/b2 = c1/c2 for?

OpenStudy (anonymous):

nvm I got it.

OpenStudy (anonymous):

We have to take the equations in the form a1x+b1y+c1=0;a2x+b2y+c2=0;

OpenStudy (anonymous):

\[\bf \frac{ -3 }{ b-2 }=\frac{ 5 }{ 3 } \implies b-2=\frac{ -9 }{ 5 } \implies b = \frac{ -9 }{ 5 }+2=\frac{ 1 }{ }\]\[\bf \frac{ 2a-1 }{ 3 }=\frac{ 5 }{ 3 } \implies 2a-1=5 \implies a = \frac{ 5 +1}{ 2 }=3\]

OpenStudy (anonymous):

b=1/5, the 5 doesn't show up for some reason lol.

OpenStudy (anonymous):

@bajji

OpenStudy (anonymous):

@genius12 Thanks a lot!! :D